OpenStudy (anonymous):

WILL MEDAL!! which ordered pairs in the form (x, y) are solutions to the equation 3x – 4y = 21? Choose all answers that are correct. A.(−3, 3) B.(−1, −6) C.(7, 0) D.(11, 3)

OpenStudy (anonymous):

Input the answers into the equation. A= -9-12=21 Wrong, because -9-12 is equal to -21, not 21 B=-3+24=21 Correct C=21-0=21 Correct D=33-12=21 Correct

OpenStudy (anonymous):

@redheadangel ?

OpenStudy (redheadangel):

Mateosa is correct :)

OpenStudy (anonymous):

k thx and could u help me with a few more?

OpenStudy (redheadangel):

yup :)

OpenStudy (anonymous):

Determine whether the ordered pairs in the form (x, y) are solutions of the given equation and select the correct answer. 3x - y/4 = 11 (3, -8), (4,4) A. Yes, both are solutions B. No, neither is a solution. C. The first is a solution, but the second is not. D. The first is not a solution, but the second is.

OpenStudy (redheadangel):

D :)

OpenStudy (redheadangel):

first is: 9-2=7 Second: 12-1=11

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation (x + 3)y = 14? Choose all answers that are correct: A. (−5, −7) B. (−1, 7) C. (10, −2) D. (11, 1) I think b is right @redheadangel

OpenStudy (redheadangel):

B and D are both correct

OpenStudy (redheadangel):

May I ask what grade you are in?

OpenStudy (anonymous):

u sure? im in 8th

OpenStudy (redheadangel):

yes, i am sure, and cool.

OpenStudy (anonymous):

k I just submitted and got a 60%

OpenStudy (anonymous):

B,A, AND D

OpenStudy (redheadangel):

A really?? shoot, so sorry must have calculated wrong :(